### v1

Abstract
Let $R$ be a commutative ring with identity and $S\subseteq R$ be a multiplicative set. An ideal $Q$ of $R$ (disjoint from $S$) is said to be $S$-primary if there exists an $s\in S$ such that for all $x,y\in R$ with $xy\in Q$, we have $sx\in Q$ or $sy\in rad(Q)$. Also, we say that an ideal of $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as finite intersection of $S$-primary ideals. In this paper, first we provide an example of $S$-Noetherian ring in which an ideal does not have a primary decomposition. Then our main aim of this paper is to establish the existence and uniqueness of $S$-primary decomposition in $S$-Noetherian rings as an extension of a historical theorem of Lasker-Noether.
  Keywords: $S$-Noetherian ring, $S$-primary ideal, $S$-irreducible ideal, S-primary decompostion.
MSC(2020): 13A15, 13B02, 13C05, 13E05
1 Introduction
The theory of Noetherian rings played an important role in the development of structure theory of commutative rings. One of the roots of this theory is the historical article [15] by Noether in 1921. Recall that a ring is called Noetherian if it satisfies ascending chain condition on ideals. In the past few decades, several generalizations of Noetherian rings have been extensively studied by many authors because of its importance (see [1], [3], [4], [5], [8] and [12]). As one of its crucial generalizations, Anderson and Dumitrescu [5] introduced $S$-Noetherian rings. A commutative ring $R$ with identity is called $S$-Noetherian, where $S\subseteq R$ is a given multiplicative set, if for each ideal $I$ of $R$, $sI\subseteq J\subseteq I$ for some $s\in S$ and some finitely generated ideal $J$. Theory of primary decomposition, considered as a generalization of the factorization of an integer $n\in\mathbb{Z}$ into a product of prime powers, initiated by Lasker-Noether [11, 15] in their abstract treatment of commutative rings. Lasker-Noether proved that in a commutative Noetherian ring, every ideal can be decomposed as an intersection, called primary decomposition, of finitely many primary ideals (popularly known as the Lasker-Noether decomposition theorem). Due to its segnificance, this theory quickly grew as one of the basic tools of commutative algebra and algebraic geometry. It gives an algebraic foundation for decomposing an algebraic variety into its irreducible components. Recently in [14], Massaoud introduced the concept of $S$-primary ideals as a proper generalization of primary ideals. Let $R$ be a commutative ring with identity and $S\subseteq R$ a multiplicative set. A proper ideal $Q$ (disjoint from $S$) of $R$ is said to be $S$-primary if there exists an $s\in S$ such that for all $a,b\in R$ if $ab\in Q$, then $sa\in Q$ or $sb\in rad(Q)$. The author [14] investigated several properties of this class of ideals and showed that $S$-primary ideals enjoy analogue of many properties of primary ideals. Given the significance of primary decomposition in Noetherian rings, a natural question arises: can we extend the concept of primary decomposition from Noetherian rings to $S$-Noetherian rings ? To extend primary decomposition from Noetherian rings to the broader class of $S$-Noetherian rings, a natural approach is to replace "primary ideals" by "$S$-primary ideals" in the decomposition. By doing this, we can logically adapt this powerful concept that allows us to extend variuos structural properties of Noetherian rings to $S$-Noetherian rings.
In this paper, we introduce the concept of $S$-primary decomposition as a generalization of primary decomposition. We say that an ideal (disjoint with $S$) of a ring $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$. First we provide an example of $S$-Noetherian ring in which primary decomposition does not exist (see Example 1) which assert that an $S$-Noetherian ring need not be a Laskerian ring in general. Then as one of our main results, we establish the existence of $S$-primary decomposition in $S$-Noetherian rings as a generalization of historical Lasker-Noether decomposition theorem (see Theorem 10 and Theorem 11). Among the other results, we extend first and second uniqueness theorems of primary decomposition to $S$-primary decomposition (see Theorem 16 and Theorem 19).
Throughout the paper, $R$ will be a commutative ring with identity and $S$ be a multiplicative set of $R$ unless otherwise stated.
2 $S$-Primary Decomposition in $S$-Noetherian Ring
It is well known that primary decomposition exists in a Noetherian ring. Recall from [1] that a ring $R$ is said to have a Noetherian spectrum if $R$ satisfies the ascending chain condition (ACC) on radical ideals. This is equivalent to the condition that $R$ satisfies the ACC on prime ideals, and each ideal has only finitely many prime ideals minimal over it. Also, a ring $R$ is said to be Laskerian if each ideal of $R$ has a primary decomposition. We begin by providing an example of an $S$-Noetherian ring in which primary decomposition does not hold.
Example 1.
Let $R=F[x_{1},x_{2},\ldots,x_{n},\ldots]$ be a polynomial ring in infinitely many indeterminates over a field $F$. Since $R$ has an ascending chain of prime ideals $(x_{1})\subseteq(x_{1},x_{2})\subseteq\cdots\subseteq(x_{1},x_{2},\ldots,x_{n})\subseteq\cdots$ which does not terminate, so $R$ has no Noetherian spectrum. This implies that $R$ is a non-Laskerian ring since every Laskerian ring has a Noetherian spectrum (see [10, Theorem 4]). Consider a multiplicative set $S=R\setminus\{0\}$. Then by [5, Proposition 2(a)], $R$ is an $S$-Noetherian ring. Hence $R$ is an $S$-Noetherian ring but not Laskerian.
Thus we need $S$-version of primary decomposition of ideals. Now we define the concept of $S$-primary decomposition of ideals as a generalization of primary decomposition.
Definition 2.
An ideal $Q$ (disjoint with $S$) of a ring $R$ is said to be $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$. Further, an $S$-primary decomposition $I=\bigcap\limits_{i=1}^{n}Q_{i}$ is said to be minimal if the $rad(Q_{i})$ are all distinct, and $\bigcap\limits_{j\neq i}Q_{j}\nsubseteq Q_{i}$ for all $i=1,2,\ldots,n$.
It is clear from the definition that the concepts of $S$-primary decomposition and primary decomposition coincide for $S=\{1\}$. The following examples show that the concept of $S$-primary decomposition is a proper generalization of the concept of primary decomposition.
Example 3.
[…]
Example 4.
Consider the Boolean ring $R=\mathcal{P}(\mathbb{N})$, the power set of $\mathbb{N}$, where addition in $R$ being the symmetric difference of sets, and multiplication being the intersection of sets. By [17, Theorem 1], the zero ideal $(0)$ in $R$ does not have a primary decomposition. Clearly, $R$ is not Noetherian. Consider a multiplicative set $S=R\setminus\{0\}$. Then by [5, Proposition 2(a)], $R$ is an $S$-Noetherian ring. Next, we show that $(0)$ is an $S$-primary ideal of $R$. For this, let $A,B\in R$ such that $AB=0$ and $sA\neq 0$ for all $s\in S$. This implies that $A$ is a non-zero element of $R$, and hence $A\in S$. Put $s=A$. Then we have $sB=0$, as desired. Hence $(0)$ is $S$-primary decomposable.
Recall from [7], an ideal $I$ of the ring $R$ is called irreducible if $I=J\cap K$ for some ideals $J$, $K$ of $R$, then either $I=J$ or $I=K$. It is well known that the classical proof of existence of primary decomposition in Noetherian ring involves the concept of irreducible ideals. So we need $S$-version of irreducible ideals to prove the existence of $S$-primary decomposition in $S$-Noetherian rings.
Definition 5.
An ideal $Q$ (disjoint with $S$) of the ring $R$ is called $S$-irreducible if $s(I\cap J)\subseteq Q\subseteq I\cap J$ for some $s\in S$ and some ideals $I$, $J$ of $R$, then there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$.
It is clear from the definition that every irreducible ideal is an $S$-irreducible ideal. However, the following example shows that an $S$-irreducibile ideal need not be irreducible.
Example 6.
Let $R=F[x,y,z]$, where $F$ is a field, and $S=\{z^{k}\hskip 2.84544pt|\hskip 2.84544ptk\in\mathbb{N}\cup\{0\}\}$ be a multiplicative set. Since $R$ is a Noetherian ring, so $R$ is an $S$-Noetherian ring. Then the ideal $Q=(x,yz)$ is not irreducible since $Q=(x,y)\cap(x,z)$. On the other hand if we take $s=z\in S$, then $s(x,y)=(sx,sy)\subseteq Q$. Thus $Q$ is an $S$-irreducible ideal of $R$.
Recall from [2], an ideal $Q$ (disjoint with $S$) of a ring $R$ is said to be an $S$-prime if there exists an $s\in S$ such that for $a,b\in R$ with $ab\in Q$, we have either $sa\in Q$ or $sb\in Q$. Clearly every $S$-prime ideal is $S$-primary. The following example shows that converse of this is not true in general.
Example 7.
Consider $R=\mathbb{Z}$, $Q=4\mathbb{Z}$, and $S=\mathbb{Z}\setminus 2\mathbb{Z}$. Notice that $4\in Q$ but $2s\notin Q$ for all $s\in S$. This implies that $Q$ is not an $S$-prime ideal of $R$. Now we prove that $Q$ is an $S$-primary ideal of $R$. For this, let $a,b\in R$ such that $ab\in Q$. Then we have two cases:
Case-1 If $a=2r$ and $b=2k$, where $r$, $k\in\mathbb{Z}$, then $ab\in Q$. In this case, there always exists $s\in S$ such that either $sa\in rad(Q)$ or $sb\in rad(Q)$. Thus $Q$ is $S$-primary.
Case-2 If either $a=4m$ or $b=4n$, where $m$, $n\in\mathbb{Z}$, then $ab\in Q$. Also in this case, there always exists $s\in S$ such that either $sa\in rad(Q)$ or $sb\in rad(Q)$. Therefore $Q$ is a $S$-primary ideal of $R$.
Recall from [14] that if $Q$ is an $S$-primary ideal of a ring $R$ and $rad(Q)=P$ is an $S$-prime ideal, then $Q$ is said to be an $S$-$P$-primary ideal of $R$.
Proposition 8.
Finite intersection of $S$-$P$-primary ideals is $S$-$P$-primary.
Proof.
Let $Q_{1},Q_{2},\ldots,Q_{n}$ be $S$-$P$-primary ideals, then $S\cap Q_{i}=\emptyset$ for each $i=1,2,\ldots,n$, and so $S\cap(\bigcap\limits_{i=1}^{n}Q_{i})=\emptyset$. Suppose $Q=\bigcap\limits_{i=1}^{n}Q_{i}$. Since each $Q_{i}$ is $S$-$P$-primary, $rad(Q)=rad(\bigcap\limits_{i=1}^{n}Q_{i})=\bigcap\limits_{i=1}^{n}rad(Q_{i})=P$. Now let $xy\in Q$ with $sy\notin Q$ for all $s\in S$. Consequently, there exists $k$, $(1\leq k\leq n)$ such that $xy\in Q_{k}$ and $sy\notin Q_{k}$ for all $s\in S$. This implies that there exists $s\in S$ such that $sx\in rad({Q_{k}})=P$. Hence $Q$ is $S$-$P$-primary. ∎
In view of the Proposition 8, we can always obtain minimal $S$-primary decomosition from $S$-primary decomosition by replacing all $S$-primary ideals with the same radical by their intersection. Following [6], let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $\mathcal{I}$ (disjoint with $S$) be an ideal of $R$. Then $\mathcal{I}$ is said to be $S$-maximal if there exists $s\in S$ such that whenever $\mathcal{I}\subseteq J$ for some ideal $J$ of $R$, then either $sJ\subseteq\mathcal{I}$ or $J\cap S\neq\emptyset$.
Proposition 9.
Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ (disjoint with $S$) be an ideal of $R$. If $rad(Q)=\mathcal{M}$ is $S$-maximal, then $Q$ is an $S$-primary ideal of $R$.
Proof.
Let $a,b\in R$ such that $ab\in Q$ and $sb\notin rad{(Q)}=\mathcal{M}$ for all $s\in S$. Choose a maximal ideal $\mathcal{M^{\prime}}$ of $R$ such that $\mathcal{M}\subseteq\mathcal{M^{\prime}}$ and $\mathcal{M^{\prime}}\cap S=\emptyset$. This implies that there exists $s^{\prime}\in S$ such that $s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}$ since $\mathcal{M}$ is $S$-maximal. Consequently, $sb\notin s^{\prime}\mathcal{M^{\prime}}$ since $sb\notin\mathcal{M}$ for all $s\in S$. If $b\in\mathcal{M^{\prime}}$, then $s^{\prime}b\in s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}$, a contradiction since $sb\notin\mathcal{M}$ for all $s\in S$. Therefore $b\notin\mathcal{M^{\prime}}$. This implies that $\mathcal{M^{\prime}}+Rb=R$. Write $1=c+\lambda b$ for some $c\in\mathcal{M^{\prime}}$ and $\lambda\in R$, then $s^{\prime}=s^{\prime}c+\lambda s^{\prime}b$. Now since $s^{\prime}c\in s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}=rad(Q)$, there exists $k\geq 1$ such that $(s^{\prime}c)^{k}\in Q$. Then by binomial theorem $(s^{\prime})^{k}=(s^{\prime}c+\lambda s^{\prime}b)^{k}=\binom{k}{0}(s^{\prime}c)^{k}+\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda s^{\prime}b)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda s^{\prime}b)^{2}+\cdots+\binom{k}{k}(s^{\prime}c)^{k-k}(\lambda s^{\prime}b)^{k}=(s^{\prime}c)^{k}+s^{\prime}b\left(\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda)^{2}(s^{\prime}b)+\cdots+(\lambda)^{k}(s^{\prime}b)^{k-1}\right)$=$(s^{\prime}c)^{k}+\lambda^{\prime}s^{\prime}b$ where $\lambda^{\prime}=\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda)^{2}(s^{\prime}b)+\cdots+(\lambda)^{k}(s^{\prime}b)^{k-1}\in R$. Then $(s^{\prime})^{k}a=a(s^{\prime}c)^{k}+\lambda^{\prime}s^{\prime}ab$, and so $(s^{\prime})^{k}a\in Q$ since $(s^{\prime}c)^{k}\in Q$ and $ab\in Q$. Consequently, $(s^{\prime}a)^{k}=(s^{\prime})^{k}a(a)^{k-1}\in Q$ since $(s^{\prime})^{k}a\in Q$. Hence $Q$ is an $S$-primary ideal of $R$. ∎
Following [3], let $E$ be a family of ideals of a ring $R$. An element $I\in E$ is said to be an $S$-maximal element of $E$ if there exists an $s\in S$ such that for each $J\in E$, if $I\subseteq J$, then $sJ\subseteq I$. Also an increasing sequence $(I_{i})_{i\in\mathbb{N}}$ of ideals of $R$ is called $S$-stationary if there exists a positive integer $k$ and $s\in S$ such that $sI_{n}\subseteq I_{k}$ for all $n\geq k$.
The following theorem provides a connection between the concepts of $S$-irreducible ideals and $S$-primary ideals.
Theorem 10.
Let R be an $S$-Noetherian ring. Then every $S$-irreducible ideal of $R$ is $S$-primary.
Proof.
Suppose $Q$ is an $S$-irreducible ideal of $R$. Let $a,b\in R$ be such that $ab\in Q$ and $sb\notin Q$ for all $s\in S$. Our aim is to show that there exists $t\in S$ such that $ta\in rad(Q).$ Consider $A_{n}=\{x\in R\hskip 5.69046pt|\hskip 2.84544pta^{n}x\in Q\}$ for $n\in\mathbb{N}.$ Clearly each $A_{n}$ is an ideal of $R$ and $A_{1}\subseteq A_{2}\subseteq A_{3}\subseteq\cdots$ is an increasing chain of ideals of $R$. Since $R$ is $S$-Noetherian, then by [9, Theorem 2.3], this chain is $S$-stationary, i.e., there exist $k\in\mathbb{N}$ and $s\in S$ such that $sA_{n}\subseteq A_{k}$ for all $n\geq k$. Consider the two ideals $I=\textless a^{k}\textgreater+\hskip 2.84544ptQ$ and $J=\textless b\textgreater+\hskip 2.84544ptQ$ of $R$. Then $Q\subseteq I\cap J$. For the reverse containment, let $y\in I\cap J$. Write $y=a^{k}z+q$ for some $z\in R$ and $q\in Q$. Since $ab\in Q$, $aJ\subseteq Q$; whence $ay\in Q$. Now $a^{k+1}z=a(a^{k}z)=a(y-q)\in Q$. This implies that $z\in A_{k+1}$, and so $sz\in sA_{k+1}\subseteq A_{k}$. Consequently, $a^{k}sz\in Q$ which implies that $a^{k}sz+sq=sy\in Q$. Thus we have $s(I\cap J)\subseteq Q\subseteq I\cap J$. This implies that there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$ since $Q$ is $S$-irreducible. If $ss^{\prime}J\subseteq Q$, then $ss^{\prime}b\in Q$ which is not possible. Therefore $ss^{\prime}I\subseteq Q$ which implies that $ss^{\prime}a^{k}\in Q$. Put $t=ss^{\prime}\in S$. Then $(ta)^{k}\in Q$ and hence $ta\in rad(Q)$, as desired. ∎
Now we are in a position to prove the existence of $S$-primary decomposition in $S$-Noetherian rings as our main result.
Theorem 11.
(Existence of $S$-Primary Decomposition) Let $R$ be an $S$-Noetherian ring. Then every proper ideal (disjoint with $S$) of $R$ can be written as a finite intersection of $S$-primary ideals.
Proof.
Let $E$ be a collection of ideals of $R$ which can not be written as a finite intersection of $S$-irreducible ideals. We wish to show $E=\emptyset$. On contrary suppose $E\neq\emptyset$. Since $R$ is an $S$-Noetherian ring, by [9, Theorem 2.3], there exists an $S$-maximal element in $E$, say $I$. Evidently, $I$ is not an $S$-irreducible ideal, and so $I$ is not an irreducible ideal. This implies that $I=J\cap K$ for some ideals $J$ and $K$ of $R$ with $I\neq J$ and $I\neq K$. Since $I$ is not $S$-irreducible, $sJ\nsubseteq I$ and $sK\nsubseteq I$ for all $s\in S$. Now we claim that $J$, $K\notin E$. For this, suppose that $J$ and $K$ are in $E$, then there exist $s^{\prime},s^{\prime\prime}\in S$ such that $s^{\prime}J\subseteq I$ and $s^{\prime\prime}K\subseteq I$ and $I\subset J$, $I\subset K$ which is not possible. Therefore $J,K\notin E$. This implies that $J$ and $K$ can be written as a finite intersections of $S$-irreducible ideals. Consequently, $I$ can also be written as a finite intersection of $S$-irreducible ideals since $I=J\cap K$, a contradiction as $I\in E$. Thus $E=\emptyset$, i.e., every proper ideal of $R$ can be written as finite intersection of $S$-irreducible ideals. But by Theorem 10, an $S$-irreducible ideal is $S$-primary. Therefore every ideal of $R$ can be written as a finite intersection of $S$-primary ideals. ∎
Recall that if $I$ is any ideal of $R$, the radical of $I$ is $rad(I)=\{x\in R\hskip 2.84544pt|\hskip 2.84544ptx^{n}\in I\hskip 2.84544ptfor\hskip 2.84544ptsome\hskip 2.84544ptn\geq 0\}$. If $I=rad(I)$, then $I$ is called radical ideal.
As an application of this theorem, we provide the following result.
Corollary 12.
Let $R$ be an $S$-Noetherian ring. Then every radical ideal $I$ (disjoint with $S$) of $R$ is the intersection of finitely many $S$-prime ideals.
Proof.
By Theorem 11, there exist finitely many $S$-primary ideals $Q_{1},Q_{2},\ldots,Q_{r}$ such that $I=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$. Also by [14, Proposition 2.5], $rad({Q_{i}})=P_{i}$ is $S$-prime for $i=1,2,\ldots,r$. Consequently, $I=rad(I)=P_{1}\cap P_{2}\cap\cdots\cap P_{r}$. This completes the proof. ∎
Remark 13.
Let $R$ be an $S$-Noetherian ring and $I$ be an ideal of $R$. Then by Theorem 11, $I$ has $S$-primary decomposition. Suppose $I=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$ is an $S$-primary decomposition of $I$. Since each $Q_{i}$ is $S$-primary, so by [16, Proposition 2.6], each $S^{-1}Q_{i}$ is primary. Consequently, $S^{-1}I=S^{-1}Q_{1}\cap S^{-1}Q_{2}\cap\cdots\cap S^{-1}Q_{r}$ is a primary decomposition of $S^{-1}I$
To prove $S$-version of the $1st$ uniqueness theorem for $S$-Noetherian rings, we need the following lemmas:
Lemma 14.
Let $Q_{1},Q_{2},\ldots,Q_{n}$ be ideals of $R$, and $P$ be an $S$-prime ideal containing $\bigcap\limits_{k=1}^{n}Q_{k}$. Then there exists $s\in S$ such that $sQ_{k}\subseteq P$ for some $k$. In particular, if $P=\bigcap\limits_{k=1}^{n}Q_{k}$, then $sQ_{k}\subseteq P\subseteq Q_{k}$ for some $k$.
Proof.
On contrary suppose $sQ_{k}\nsubseteq P$ for all $s\in S$ and $k=1,2,\dots,n$. Then for each $k$, there exists $x_{k}\in Q_{k}$ such that $sx_{k}\notin P$ for all $s\in S$. Evidently, $x_{1}x_{2}\ldots x_{n}\in Q_{1}Q_{2}\ldots Q_{n}$. This implies that $(sx_{1})(sx_{2})\cdots(sx_{n})\in Q_{1}Q_{2}\ldots Q_{n}\subseteq\bigcap\limits_{k=1}^{n}Q_{k}\subseteq P$. Then there exist $s^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $s^{\prime}sx_{j}\in P$ since $P$ is an $S$-prime ideal, a contradiction. Hence there exist $s\in S$ and $k\in\{1,2,\ldots,n\}$ such that $sQ_{k}\subseteq P$. Finally, if $P=\bigcap\limits_{k=1}^{n}Q_{k}$, then by above argument there exist $s\in S$ and $k\in\{1,2,\ldots,n\}$ such that $sQ_{k}\subseteq P\subseteq Q_{k}$. ∎
Lemma 15.
Let $Q$ be an $S$-primary ideal and $x$ an element of $R$ such that $sx\notin Q$ for all $s\in S$. Then $(Q:sx)$ is $S$-primary, and there exists $t\in S$ such that $t(rad(Q:sx))\subseteq rad(Q)\subseteq rad(Q:sx)$ for all $s\in S$.
Proof.
Suppose $(Q:sx)\cap S\neq\emptyset$ for some $s\in S$. Then there exists $t\in S$ such that $t\in(Q:sx)$ which imples that $tsx\in Q$, a contradiction. Thus $(Q:sx)\cap S=\emptyset$ for all $s\in S.$ Now let $s\in S$ and $yz\in(Q:sx)$ be such that $s^{\prime}y\notin rad(Q:sx)$ for all $s^{\prime}\in S$. Then $sxyz\in Q$. Since $Q$ is $S$-primary and $s^{\prime}y\notin rad(Q:sx)$ for all $s^{\prime}\in S$, there exists $s^{\prime\prime}\in S$ such that $s^{\prime\prime}sxz\in Q$. Consequently, $s^{\prime\prime}z\in(Q:sx)$. Thus $(Q:sx)$ is $S$-primary. Further, let $w\in rad(Q:sx)$. Then $w^{n}\in(Q:sx)$ for some $n\in\mathbb{N}$. This implies that $sxw^{n}\in Q$. Since $Q$ is an $S$-primary, there exists $t\in S$ such that $(tw^{n})^{k}\in Q$ for some $k$ as $tsx\notin Q$. This concludes that $(tw)^{nk}=t^{nk-k}(tw^{n})^{k}\in Q$ which implies that $tw\in rad(Q)$. Thus $t(rad(Q:sx))\subseteq rad(Q)$. Also $Q\subseteq(Q:sx)$ implies that $rad(Q)\subseteq rad(Q:sx)$, and hence $t(rad(Q:sx))\subseteq rad(Q)\subseteq rad(Q:sx)$. ∎
Theorem 16.
[…]
Proof.
For any $x\in R$, we have $(I:x)=(\bigcap\limits_{i=1}^{n}Q_{i}:x)=\bigcap\limits_{i=1}^{n}(Q_{i}:x)$, and so $rad(I:x)=\bigcap\limits_{i=1}^{n}rad(Q_{i}:x)$. First we show that $rad(I:x)$ satisfies $sP_{i}\subseteq s^{\prime}(rad(I:x))\subseteq P_{i}$ for some $s,s^{\prime}\in S$ is $S$-prime. For this, let $\alpha\beta\in rad(I:x)$ such that $t\beta\notin rad(I:x)$ for all $t\in S$. This implies $s^{\prime}\alpha\beta\in s^{\prime}rad(I:x)\subseteq P_{i}$. Then there exists $s^{\prime\prime}\in S$ such that either $s^{\prime\prime}\alpha\in P_{i}$ or $s^{\prime}s^{\prime\prime}\beta\in P_{i}$ since $P_{i}$ is $S$-prime. If $s^{\prime}s^{\prime\prime}\beta\in P_{i}$, then $ss^{\prime}s^{\prime\prime}\beta\in sP_{i}\subseteq s^{\prime}(rad(I:x))\subseteq rad(I:x)$, a contradiction since $t\beta\notin rad(I:x)$ for all $t\in S$. Hence $s^{\prime\prime}\alpha\in P_{i}$, and so $ss^{\prime\prime}\alpha\in sP_{i}\subseteq s^{\prime}rad(I:x)\subseteq rad(I:x)$, i.e., $s^{\prime\prime\prime}\alpha\in rad(I:x)$, where $s^{\prime\prime\prime}=ss^{\prime\prime}\in S$. Thus $rad(I:x)$ is an $S$-prime ideal. Now by Lemma 14, there exist $t^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $t^{\prime}(rad(Q_{j}:x))\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$. Choose $x$ such that $tx\notin Q_{j}$ for all $t\in S$. Then by Lemma 15, there exists $t^{\prime\prime}\in S$ such that $t^{\prime\prime}(rad(Q_{j}:x))\subseteq rad(Q_{j})=P_{j}\subseteq rad(Q_{j}:x)$. This implies that $t^{\prime}t^{\prime\prime}(rad(Q_{j}:x))\subseteq t^{\prime}P_{j}\subseteq t^{\prime}(rad(Q_{j}:x))\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$. Thus we have $t^{\prime}P_{j}\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$, and so $t^{\prime\prime}t^{\prime}P_{j}\subseteq t^{\prime\prime}(rad(I:x))\subseteq t^{\prime\prime}(rad(Q_{j}:x))\subseteq rad(Q_{j})=P_{j}$. Put $s=t^{\prime\prime}t^{\prime}$ and $s^{\prime}=t^{\prime\prime}$. Then we have $sP_{j}\subseteq s^{\prime}(rad(I:x))\subseteq P_{j}$ for some $j$. Conversely, since decomposition is minimal, so for each $k$, $\bigcap\limits_{i\neq k}^{n}Q_{i}\nsubseteq Q_{k}$. This implies that there exists $x_{k}\in(\bigcap\limits_{i\neq k}^{n}Q_{i})\setminus Q_{k}$ for each $k$. Consequently, for each $k$, we have $(I:x_{k})=\bigcap\limits_{i=1}^{n}(Q_{i}:x_{k})=(Q_{k}:x_{k})$ since $(Q_{i}:x_{k})=R$ for all $i\neq k$ by [7, Lemma 4.4]. Notice that $tx_{k}\notin Q_{k}$ for every $t\in S$, for if $tx_{k}\in Q_{k}$ for some $t\in S$, then $t\in(Q_{k}:x_{k})=(I:x_{k})\subseteq rad(I:x_{k})$, which is not possible since $rad(I:x_{k})$ is $S$-prime. So by Lemma 15, for each $k$, there exists $s_{k}\in S$ such that $s_{k}(rad(Q_{k}:x_{k}))\subseteq rad(Q_{k})=P_{k}$. Also $Q_{k}\subseteq(Q_{k}:x_{k})=(I:x_{k})$, so $P_{k}=rad(Q_{k})\subseteq rad(I:x_{k})$. Consequently, for each $k$, $s_{k}P_{k}\subseteq s_{k}(rad(I:x_{k}))=s_{k}(rad(Q_{k}:x_{k}))\subseteq P_{k}$. Put $s=s_{1}s_{2}\ldots s_{n}$. Then $sP_{k}\subseteq s(rad(I:x_{k}))\subseteq P_{k}$ for all $k=1,2,\ldots,n$. This completes the proof. ∎
For any ideal $I$ of a ring $R$ and any multiplicative set $S\subseteq R$, the contraction of ideal $S^{-1}I$ is denoted by $S(I)=\{a\in R\hskip 5.69046pt|\hskip 5.69046pt\dfrac{a}{1}\in S^{-1}I\}$. Notice that $I\subseteq S(I)$. To prove the second uniqueness theorem for the $S$-Noetherian rings, we need the following results:
Lemma 17.
Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ be an $S$-$P$-primary ideal of $R$. If $S\cap P=\emptyset$, then $S^{-1}Q$ is $S^{-1}P$-primary and its contraction in $R$ satisfies $tS(Q)\subseteq Q\subseteq S(Q)$ for some $t\in S$.
Proof.
Clearly $S^{-1}Q\neq S^{-1}R$, for $S\cap P=\emptyset$ and $P=rad(Q)$. Let $x\in S(Q)$. Then $\dfrac{x}{1}=\dfrac{a}{s}$ for some $a\in Q$ and $s\in S$. Consequently, there exists $s^{\prime}\in S$ such that $(xs-a)s^{\prime}=0$; whence $ss^{\prime}x=s^{\prime}a\in Q$. Put $t=ss^{\prime}$. Then $tx\in Q$, and so $tS(Q)\subseteq Q\subseteq S(Q)$. Now since $Q$ is $S$-primary, by [16, Proposition 2.6], $S^{-1}Q$ is a primary ideal of $S^{-1}R$ and $rad(S^{-1}Q)=S^{-1}(rad(Q))=S^{-1}P$. Also since $P$ is an $S$-prime ideal of $R$, by [2, Remark $1$], $S^{-1}P$ is a prime ideal of $S^{-1}R$. Thus $S^{-1}Q$ is an $S^{-1}P$-primary ideal of $S^{-1}R$.
Lemma 18.
Let $R$ be an $S$-Noetherian ring, and $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $S$-$P_{i}$-primary. If $S\cap\left(\bigcap\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ for some $m$. Then $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$ and $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S(I)$ for some $t\in S$.
Proof.
By [7, Proposition 3.11(v)], $S^{-1}I=\bigcap\limits_{i=1}^{n}S^{-1}Q_{i}$. Since $S\cap Q_{i}\neq\emptyset$ for $i=m+1,m+2,\ldots,n$, so by [7, Proposition 4.8(i)], $S^{-1}Q_{i}=S^{-1}R$ for $i=m+1,m+2,\ldots,n$. Consequently, $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$. Also by Lemma 17, $S^{-1}Q_{i}$ is $S^{-1}P_{i}$-primary for $i=1,2,\ldots,m$. Since $P_{i}$ are distinct, $S^{-1}P_{i}$ are also distinct, and so $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$ is a minimal $S$-primary decomposition. Next, let $x\in S(I)$. Then $\dfrac{x}{1}\in S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$. Write $\dfrac{x}{1}=\dfrac{a_{i}}{s_{i}}$ for some $a_{i}\in Q_{i}$ and $s_{i}\in S$. This implies that $(xs_{i}-a_{i})s=0$ for some $s\in S$; whence $s_{i}sx=sa_{i}\in Q_{i}$ for $i=1,2,\ldots,m$. Put $t=ss_{1}s_{2}\ldots s_{m}$. Then $tx\in\bigcap\limits_{i=1}^{m}Q_{i}$, and hence $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}$. Also by [7, Proposition 1.17(i)], $\bigcap\limits_{i=1}^{m}Q_{i}\subseteq\bigcap\limits_{i=1}^{m}S(Q_{i})=S(I)$. Hence $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S(I)$. ∎
For an ideal $I$ of $R$, let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $S$-$P_{i}$-primary. Then the $S$-prime ideals $P_{i}$ are said to belong to $I$. Also we say that $P_{i}$ ($1\leq i\leq n$) is isolated $S$-prime if $sP_{j}\nsubseteq P_{i}$ for all $s\in S$ and for all $j\neq i$. Other $S$-prime ideals are called embedded $S$-prime ideals.
Theorem 19.
($S$-version of 2nd uniqueness theorem). Let $R$ be an $S$-Noetherian ring and $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition of $I$, where $Q_{i}$ is $S$-$P_{i}$-primary. If $\{P_{{1}},\ldots,P_{{m}}\}$ is a set of isolated $S$-prime ideals of $I$ for some $m$ $(1\leq m\leq n)$, then $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition.
Proof.
Since each $P_{i}$ is $S$-prime, so by [2, Proposition 1], there exist $s_{i}\in S$ such that $(P_{i}:s_{i})$ is a prime ideal for $i=1,2,\ldots,m$. Consider a multiplicative set $S^{\prime}=R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$ of $R$. Evidently, $S^{\prime}\cap P_{i}=\emptyset$ for $i=1,2,\ldots m$ since $P_{i}\subseteq(P_{i}:s_{i})$. We claim that $S^{\prime}\cap P_{k}\neq\emptyset$ for $k=m+1,m+2,\ldots,n$. On contary $S^{\prime}\cap P_{k}=\emptyset$ for some $k$. This implies that $P_{k}\subseteq\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$, so by [7, Proposition 1.11(i)], $P_{k}\subseteq(P_{j}:s_{j})$ for some $j$ $(1\leq j\leq m)$. Consequently, $s_{j}P_{k}\subseteq P_{j}$, a contradiction since $P_{j}$ is isolated $S$-prime. This concludes that $S^{\prime}\cap\left(\bigcap\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S^{\prime}\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ since $P_{i}=rad(Q_{i})$ for all $i=1,2,\ldots,n.$ Now since each $P_{i}$ is $S$-prime, $S\cap P_{i}=\emptyset$ for all $i=1,2,\ldots,m$. This implies that $S\cap(P_{i}:s_{i})=\emptyset$ for all $i=1,2,\ldots,m$. For, if $S\cap(P_{i}:s_{i})\neq\emptyset$ for some $i$, then there exits $s^{\prime}\in S$ such that $s^{\prime}s_{i}\in P_{i}$, a contradiction. Consequently, $S\cap(\bigcup\limits_{i=1}^{m}(P_{i}:s_{i}))=\emptyset$, and so $S\subseteq R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})=S^{\prime}$. Then by [5, Proposition 2(d)], $R$ is $S^{\prime}$-Noetherian. Also since each $Q_{i}$ is $S$-primary and $S^{\prime}\cap Q_{i}=\emptyset$ for all $i$ $(1\leq i\leq m)$, so $Q_{i}$ is $S^{\prime}$-primary for all $i=1,2,\ldots,m.$ Thus by Lemma 18, we have $tS(I)\subseteq\bigcap\limits_{j=1}^{m}Q_{j}\subseteq S(I)$ for some $t\in S$. Hence $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition. ∎

### v2

@@
-
+A Study of $S$-Primary Decompositions
 Abstract
-Let $R$ be a commutative ring with identity and $S\subseteq R$ be a multiplicative set.
+Let $R$ be a commutative ring with identity, and let $S\subseteq R$ be a multiplicative set.
 An ideal $Q$ of $R$ (disjoint from $S$) is said to be $S$-primary if there exists an $s\in S$ such that for all $x,y\in R$ with $xy\in Q$, we have $sx\in Q$ or $sy\in rad(Q)$.
-Also, we say that an ideal of $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as finite intersection of $S$-primary ideals.
-In this paper, first we provide an example of $S$-Noetherian ring in which an ideal does not have a primary decomposition.
+Also, we say that an ideal of $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals.
+In this paper, first we provide an example of an $S$-Noetherian ring in which an ideal does not have a primary decomposition.
 Then our main aim of this paper is to establish the existence and uniqueness of $S$-primary decomposition in $S$-Noetherian rings as an extension of a historical theorem of Lasker-Noether.
-  Keywords: $S$-Noetherian ring, $S$-primary ideal, $S$-irreducible ideal, S-primary decompostion.
-MSC(2020): 13A15, 13B02, 13C05, 13E05
+Keywords: $S$-Noetherian ring, $S$-primary ideal, $S$-irreducible ideal, S-primary decomposition.
+MSC(2010): 13C05, 13B02, 13E05.
 1 Introduction
-The theory of Noetherian rings played an important role in the development of structure theory of commutative rings.
+The theory of Noetherian rings has been playing an important role in the development of structure theory of commutative rings.
 One of the roots of this theory is the historical article [14] by Noether in 1921.
 Recall that a ring is called Noetherian if it satisfies ascending chain condition on ideals.
-In the past few decades, several generalizations of Noetherian rings have been extensively studied by many authors because of its importance (see [1], [3], [4], [5], [8] and [12]).
+In the past few decades, several generalizations of Noetherian rings have been extensively studied by many authors because of its importance (see [1], [3], [4], [5], [7], and [12]).
 As one of its crucial generalizations, Anderson and Dumitrescu [5] introduced $S$-Noetherian rings.
@@
 Lasker-Noether proved that in a commutative Noetherian ring, every ideal can be decomposed as an intersection, called primary decomposition, of finitely many primary ideals (popularly known as the Lasker-Noether decomposition theorem).
-Due to its segnificance, this theory quickly grew as one of the basic tools of commutative algebra and algebraic geometry.
+Due to its significance, this theory quickly grew as one of the basic tools of commutative algebra and algebraic geometry.
 It gives an algebraic foundation for decomposing an algebraic variety into its irreducible components.
@@
 The author [13] investigated several properties of this class of ideals and showed that $S$-primary ideals enjoy analogue of many properties of primary ideals.
-Given the significance of primary decomposition in Noetherian rings, a natural question arises: can we extend the concept of primary decomposition from Noetherian rings to $S$-Noetherian rings ?
-To extend primary decomposition from Noetherian rings to the broader class of $S$-Noetherian rings, a natural approach is to replace "primary ideals" by "$S$-primary ideals" in the decomposition.
-By doing this, we can logically adapt this powerful concept that allows us to extend variuos structural properties of Noetherian rings to $S$-Noetherian rings.
+Given the significance of primary decomposition in Noetherian rings, a natural question arises:
+Question 1.1.
+Can the idea of primary decomposition in Noetherian rings be extended to $S$-Noetherian rings?
+We provide a positive answer to the above question in this paper.
+A natural way to extend primary decomposition from Noetherian rings to the broader class of $S$-Noetherian rings is to replace "primary ideals" by "$S$-primary ideals" in the decomposition process.
+By doing this, we can logically adapt this powerful concept that allows us to extend various structural properties of Noetherian rings to $S$-Noetherian rings.
 In this paper, we introduce the concept of $S$-primary decomposition as a generalization of primary decomposition.
-We say that an ideal (disjoint with $S$) of a ring $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$.
-First we provide an example of $S$-Noetherian ring in which primary decomposition does not exist (see Example 1) which assert that an $S$-Noetherian ring need not be a Laskerian ring in general.
+We say that an ideal (disjoint from $S$) of a ring $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$.
+First we provide an example of an $S$-Noetherian ring in which primary decomposition does not exist (see Example 1) which asserts that an $S$-Noetherian ring need not be a Laskerian ring in general.
 Then as one of our main results, we establish the existence of $S$-primary decomposition in $S$-Noetherian rings as a generalization of historical Lasker-Noether decomposition theorem (see Theorem 8 and Theorem 9).
@@
 Also, a ring $R$ is said to be Laskerian if each ideal of $R$ has a primary decomposition.
+Recall that [5], an ideal $I$ of $R$ is called $S$-finte if $sI\subseteq J\subseteq I$ for some finitely generated ideal $J$ of $R$ and some $s\in S$.
+Then $R$ is said to be an $S$-Noetherian ring if each ideal of $R$ is $S$-finite.
 We begin by providing an example of an $S$-Noetherian ring in which primary decomposition does not hold.
 Example 1.
-Let $R=F[x_{1},x_{2},\ldots,x_{n},\ldots]$ be a polynomial ring in infinitely many indeterminates over a field $F$.
+Let $R=F[x_{1},x_{2},\ldots,x_{n},\ldots]$ be the polynomial ring in infinitely many indeterminates over a field $F$.
 Since $R$ has an ascending chain of prime ideals $(x_{1})\subseteq(x_{1},x_{2})\subseteq\cdots\subseteq(x_{1},x_{2},\ldots,x_{n})\subseteq\cdots$ which does not terminate, so $R$ has no Noetherian spectrum.
-This implies that $R$ is a non-Laskerian ring since every Laskerian ring has a Noetherian spectrum (see [10, Theorem 4]).
-Consider a multiplicative set $S=R\setminus\{0\}$.
+This implies that $R$ is a non-Laskerian ring, by [9, Theorem 4].
+Consider the multiplicative set $S=R\setminus\{0\}$.
 Then by [5, Proposition 2(a)], $R$ is an $S$-Noetherian ring.
 Hence $R$ is an $S$-Noetherian ring but not Laskerian.
-Thus we need $S$-version of primary decomposition of ideals. Now we define the concept of $S$-primary decomposition of ideals as a generalization of primary decomposition.
+Recall that let $f:R\longrightarrow S^{-1}R$ denote the usual homomorphism of rings given by $f(r)=\frac{r}{1}$.
+For any ideal $I$ of $R$, the $f^{-1}(S^{-1}I)$ called the contraction of $I$ with respect to $S$, that is, $\{a\in R\hskip 5.69046pt|\hskip 5.69046pt\dfrac{a}{1}\in S^{-1}I\}$ is denoted by $S(I)$.
+Notice that $I\subseteq S(I)$.
+Thus we need an $S$-version of primary decomposition of ideals.
+Now we define the concept of $S$-primary decomposition of ideals as a generalization of primary decomposition.
 Definition 2.
-An ideal $Q$ (disjoint with $S$) of a ring $R$ is said to be $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$.
-Further, an $S$-primary decomposition $I=\bigcap\limits_{i=1}^{n}Q_{i}$ is said to be minimal if the $rad(Q_{i})$ are all distinct, and $\bigcap\limits_{j\neq i}Q_{j}\nsubseteq Q_{i}$ for all $i=1,2,\ldots,n$.
-It is clear from the definition that the concepts of $S$-primary decomposition and primary decomposition coincide for $S=\{1\}$.
-The following examples show that the concept of $S$-primary decomposition is a proper generalization of the concept of primary decomposition.
+Let $R$ be a ring and let $S$ be a multiplicative set of $R$.
+Let $I$ be an ideal of $R$ such that $I\cap S=\emptyset$.
+We say that $I$ admits an $S$-primary decomposition if $I$ is a finite intersection of $S$-primary ideals of $R$.
+In such a case, we say that $I$ is $S$-decomposable.
+An $S$-primary decomposition $I=\bigcap\limits_{i=1}^{n}Q_{i}$ of $I$ with $rad(Q_{i})=P_{i}$ for each $i\in\{1,2,\ldots,n\}$ is said to be minimal if the following conditions hold:
+  1. 1.
+    $S(P_{i})\neq S(P_{j})$ for all distinct $i,j\in\{1,2,\ldots,n\}$.
+  2. 2.
+    $S(Q_{i})\nsupseteq\bigcap_{j\in\{1,2,\ldots,n\}\setminus\{i\}}S(Q_{j})$ for each $i\in\{1,2,\ldots,n\}$ (equivalently,
+    $S(Q_{i})\nsupseteq\bigcap_{j\in\{1,2,\ldots,n\}\setminus\{i\}}Q_{j}$ for each $i\in\{1,2,\ldots,n\})$.
+From definition 2, the concepts of $S$-primary decomposition and primary decomposition coincide for $S=\{1\}$.
+The following example shows that the concept of $S$-primary decomposition is a proper generalization of the concept of primary decomposition.
 Example 3.
-[…]
-Example 4.
-Consider the Boolean ring $R=\mathcal{P}(\mathbb{N})$, the power set of $\mathbb{N}$, where addition in $R$ being the symmetric difference of sets, and multiplication being the intersection of sets.
-By [17, Theorem 1], the zero ideal $(0)$ in $R$ does not have a primary decomposition.
+Consider the Boolean ring $R=\prod_{n=1}^{\infty}\mathbb{Z}_{2}$ (countably infinite copies of $\mathbb{Z}_{2})$.
+According to [16, Theorem 1], the zero ideal $(0)=(0,0,0,\ldots)$ in $R$ does not have the primary decomposition.
 Clearly, $R$ is not Noetherian.
-Consider a multiplicative set $S=R\setminus\{0\}$.
-Then by [5, Proposition 2(a)], $R$ is an $S$-Noetherian ring.
+Consider the multiplicative set $S=\{{1_{R}}=(1,1,1,\ldots),s=(1,0,0,\ldots)\}$.
+Let $I$ be an ideal of $R$.
+Either $s\in I$ or $s\notin I$.
+If $s\in I$, then $sI\subseteq Rs\subseteq I$ and as $Rs$ is a finitely generated ideal of $R$, it follows that $I$ is $S$-finite.
+If $s\notin I$, then $I\subseteq(0)\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\cdots$.
+In such a case, $sI\subseteq(0)\times(0)\times(0)\times\cdots\subseteq I$.
+Hence, $I$ is $S$-finite.
+This shows that $R$ is an $S$-Noetherian ring.
 Next, we show that $(0)$ is an $S$-primary ideal of $R$.
-For this, let $A,B\in R$ such that $AB=0$ and $sA\neq 0$ for all $s\in S$.
-This implies that $A$ is a non-zero element of $R$, and hence $A\in S$.
-Put $s=A$.
-Then we have $sB=0$, as desired.
-Hence $(0)$ is $S$-primary decomposable.
+First, we observe that $(0)\cap S=\emptyset$.
+Now, let $a=(a_{n})_{n\in\mathbb{N}}$, $b=(b_{n})_{n\in\mathbb{N}}\in R$ such that $ab=0$, where each $a_{i},{b_{i}}\in\mathbb{Z}_{2}$.
+This implies that $a_{n}b_{n}=0$ for all $n\in\mathbb{N}$, in particular, $a_{1}b_{1}=0$.
+Then we have either $a_{1}=0$ or $b_{1}=0$.
+If $a_{1}=0$, then $sa=0$.
+If $b_{1}=0$, then $sb=0$.
+Thus $(0)$ is an $S$-primary ideal.
+Therefore $(0)$ is $S$-primary decomposable.
 Recall from [6], an ideal $I$ of the ring $R$ is called irreducible if $I=J\cap K$ for some ideals $J$, $K$ of $R$, then either $I=J$ or $I=K$.
-It is well known that the classical proof of existence of primary decomposition in Noetherian ring involves the concept of irreducible ideals.
+It is well known that the classical proof of existence of primary decomposition in a Noetherian ring involves the concept of irreducible ideals.
 So we need $S$-version of irreducible ideals to prove the existence of $S$-primary decomposition in $S$-Noetherian rings.
 Definition 4.
-An ideal $Q$ (disjoint with $S$) of the ring $R$ is called $S$-irreducible if $s(I\cap J)\subseteq Q\subseteq I\cap J$ for some $s\in S$ and some ideals $I$, $J$ of $R$, then there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$.
+An ideal $Q$ (disjoint from $S$) of the ring $R$ is called $S$-irreducible if $s(I\cap J)\subseteq Q\subseteq I\cap J$ for some $s\in S$ and some ideals $I$, $J$ of $R$, then there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$.
 It is clear from the definition that every irreducible ideal is an $S$-irreducible ideal. However, the following example shows that an $S$-irreducibile ideal need not be irreducible.
 Example 5.
-Let $R=F[x,y,z]$, where $F$ is a field, and $S=\{z^{k}\hskip 2.84544pt|\hskip 2.84544ptk\in\mathbb{N}\cup\{0\}\}$ be a multiplicative set.
-Since $R$ is a Noetherian ring, so $R$ is an $S$-Noetherian ring.
-Then the ideal $Q=(x,yz)$ is not irreducible since $Q=(x,y)\cap(x,z)$.
-On the other hand if we take $s=z\in S$, then $s(x,y)=(sx,sy)\subseteq Q$.
-Thus $Q$ is an $S$-irreducible ideal of $R$.
-Recall from [2], an ideal $Q$ (disjoint with $S$) of a ring $R$ is said to be an $S$-prime if there exists an $s\in S$ such that for $a,b\in R$ with $ab\in Q$, we have either $sa\in Q$ or $sb\in Q$.
-Clearly every $S$-prime ideal is $S$-primary.
-The following example shows that converse of this is not true in general.
+Let $R=\mathbb{Z}$, $S=\mathbb{Z}\setminus 3\mathbb{Z}$ and $I=6\mathbb{Z}$.
+Since $I=2\mathbb{Z}\cap 3\mathbb{Z}$, therefore $I$ is not an irreducible ideal of $R$.
+Now take $s=2\in S$.
+Then $2(3\mathbb{Z})=6\mathbb{Z}\subseteq I$.
+Thus $I$ is an $S$-irreducible ideal of $R$.
+Recall from [2], an ideal $Q$ (disjoint from $S$) of a ring $R$ is said to be $S$-prime if there exists an $s\in S$ such that for $a,b\in R$ with $ab\in Q$, we have either $sa\in Q$ or $sb\in Q$.
+Clearly, every $S$-prime ideal is $S$-primary.
+The following example shows that the converse of this is not true in general.
 Example 6.
@@
 This implies that $Q$ is not an $S$-prime ideal of $R$.
-Now we prove that $Q$ is an $S$-primary ideal of $R$.
-For this, let $a,b\in R$ such that $ab\in Q$.
-Then we have two cases:
-Case-1 If $a=2r$ and $b=2k$, where $r$, $k\in\mathbb{Z}$, then $ab\in Q$. In this case, there always exists $s\in S$ such that either $sa\in rad(Q)$ or $sb\in rad(Q)$. Thus $Q$ is $S$-primary.
-Case-2 If either $a=4m$ or $b=4n$, where $m$, $n\in\mathbb{Z}$, then $ab\in Q$.
-Also in this case, there always exists $s\in S$ such that either $sa\in rad(Q)$ or $sb\in rad(Q)$.
-Therefore $Q$ is a $S$-primary ideal of $R$.
-Recall from [14] that if $Q$ is an $S$-primary ideal of a ring $R$ and $rad(Q)=P$ is an $S$-prime ideal, then $Q$ is said to be an $S$-$P$-primary ideal of $R$.
+Obviously, $Q$ is a primary ideal of $R$ and hence $S$-primary.
+Recall from [13, Proposition 2.5] that if $Q$ is an $S$-primary ideal of a ring $R$, then $P=\text{rad}(Q)$ is an $S$-prime ideal. In such a case, we say that $Q$ is an $P$-$S$-primary ideal of $R$.
 Proposition 7.
-Finite intersection of $S$-$P$-primary ideals is $S$-$P$-primary.
+Let $S$ be a multiplicative set of a ring $R$. Then the following statements hold:
+  1. 1.
+    Finite intersection of $P$-$S$-primary ideals is $P$-$S$-primary.
+  2. 2.
+    If $Q$ is a $P$-primary ideal of $R$ with $Q\cap S=\emptyset$, then for any ideal $J$ of $R$ with $J\cap S\neq\emptyset$, $Q\cap J$ is a $(P\cap rad(J))$-$S$-primary ideal of $R$.
 Proof.
-Let $Q_{1},Q_{2},\ldots,Q_{n}$ be $S$-$P$-primary ideals, then $S\cap Q_{i}=\emptyset$ for each $i=1,2,\ldots,n$, and so $S\cap(\bigcap\limits_{i=1}^{n}Q_{i})=\emptyset$.
+  1. 1.
+    Let $Q_{1},Q_{2},\ldots,Q_{n}$ be $P$-$S$-primary ideals, then $S\cap Q_{i}=\emptyset$ for each $i=1,2,\ldots,n$, and so $S\cap(\bigcap\limits_{i=1}^{n}Q_{i})=\emptyset$.
 Suppose $Q=\bigcap\limits_{i=1}^{n}Q_{i}$.
-Since each $Q_{i}$ is $S$-$P$-primary, $rad(Q)=rad(\bigcap\limits_{i=1}^{n}Q_{i})=\bigcap\limits_{i=1}^{n}rad(Q_{i})=P$.
-Now let $xy\in Q$ with $sy\notin Q$ for all $s\in S$.
-Consequently, there exists $k$, $(1\leq k\leq n)$ such that $xy\in Q_{k}$ and $sy\notin Q_{k}$ for all $s\in S$.
-This implies that there exists $s\in S$ such that $sx\in rad({Q_{k}})=P$.
-Hence $Q$ is $S$-$P$-primary.
-∎
-In view of the Proposition 8, we can always obtain minimal $S$-primary decomosition from $S$-primary decomosition by replacing all $S$-primary ideals with the same radical by their intersection.
-Following [6], let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $\mathcal{I}$ (disjoint with $S$) be an ideal of $R$.
-Then $\mathcal{I}$ is said to be $S$-maximal if there exists $s\in S$ such that whenever $\mathcal{I}\subseteq J$ for some ideal $J$ of $R$, then either $sJ\subseteq\mathcal{I}$ or $J\cap S\neq\emptyset$.
-Proposition 9.
-Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ (disjoint with $S$) be an ideal of $R$. If $rad(Q)=\mathcal{M}$ is $S$-maximal, then $Q$ is an $S$-primary ideal of $R$.
-Proof.
-Let $a,b\in R$ such that $ab\in Q$ and $sb\notin rad{(Q)}=\mathcal{M}$ for all $s\in S$.
-Choose a maximal ideal $\mathcal{M^{\prime}}$ of $R$ such that $\mathcal{M}\subseteq\mathcal{M^{\prime}}$ and $\mathcal{M^{\prime}}\cap S=\emptyset$.
-This implies that there exists $s^{\prime}\in S$ such that $s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}$ since $\mathcal{M}$ is $S$-maximal.
-Consequently, $sb\notin s^{\prime}\mathcal{M^{\prime}}$ since $sb\notin\mathcal{M}$ for all $s\in S$.
-If $b\in\mathcal{M^{\prime}}$, then $s^{\prime}b\in s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}$, a contradiction since $sb\notin\mathcal{M}$ for all $s\in S$.
-Therefore $b\notin\mathcal{M^{\prime}}$.
-This implies that $\mathcal{M^{\prime}}+Rb=R$.
-Write $1=c+\lambda b$ for some $c\in\mathcal{M^{\prime}}$ and $\lambda\in R$, then $s^{\prime}=s^{\prime}c+\lambda s^{\prime}b$.
-Now since $s^{\prime}c\in s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}=rad(Q)$, there exists $k\geq 1$ such that $(s^{\prime}c)^{k}\in Q$.
-Then by binomial theorem $(s^{\prime})^{k}=(s^{\prime}c+\lambda s^{\prime}b)^{k}=\binom{k}{0}(s^{\prime}c)^{k}+\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda s^{\prime}b)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda s^{\prime}b)^{2}+\cdots+\binom{k}{k}(s^{\prime}c)^{k-k}(\lambda s^{\prime}b)^{k}=(s^{\prime}c)^{k}+s^{\prime}b\left(\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda)^{2}(s^{\prime}b)+\cdots+(\lambda)^{k}(s^{\prime}b)^{k-1}\right)$=$(s^{\prime}c)^{k}+\lambda^{\prime}s^{\prime}b$ where $\lambda^{\prime}=\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda)^{2}(s^{\prime}b)+\cdots+(\lambda)^{k}(s^{\prime}b)^{k-1}\in R$.
-Then $(s^{\prime})^{k}a=a(s^{\prime}c)^{k}+\lambda^{\prime}s^{\prime}ab$, and so $(s^{\prime})^{k}a\in Q$ since $(s^{\prime}c)^{k}\in Q$ and $ab\in Q$.
-Consequently, $(s^{\prime}a)^{k}=(s^{\prime})^{k}a(a)^{k-1}\in Q$ since $(s^{\prime})^{k}a\in Q$.
-Hence $Q$ is an $S$-primary ideal of $R$.
-∎
+Since each $Q_{i}$ is $P$-$S$-primary, $rad(Q)=rad(\bigcap\limits_{i=1}^{n}Q_{i})=\bigcap\limits_{i=1}^{n}rad(Q_{i})=P$.
+Now, let $xy\in Q$, where $x,y\in R$ and with $sy\notin Q$ for all $s\in S$.
+Consequently, for every $s\in S$, there exists $k_{s}$ such that $xy\in Q_{k_{s}}$ and $sy\notin Q_{k_{s}}$.
+Let $s_{i}\in S$ be the element satisfying the $S$-primary property for $Q_{i}$.
+Since we have finitely many $Q_{i}$, put $s=s_{1}s_{2}\dots s_{n}\in S$.
+Now fix $s$ and assume that $xy\in Q$ but $sy\notin Q$.
+Thus, there exists $k$ such that $xy\in Q_{k}$ and $sy\notin Q_{k}$.
+Then for $s_{k}\in S$ we obtain $s_{k}x\in rad(Q_{k})=P$ or $s_{k}y\in Q_{k}$.
+The latter case gives $sy\in Q_{k}$, a contradiction.
+Thus $sx\in rad(Q)=P$, and therefore $Q$ is $P$-$S$-primary.
+  2. 2.
+    As $Q\cap S=\emptyset$, it follows that $(Q\cap J)\cap S=\emptyset$.
+By assumption, $J\cap S\neq\emptyset$.
+Let $s\in J\cap S$.
+Let $a,b\in R$ be such that $ab\in Q\cap J$.
+Either $a\in Q$ or $b\in rad(Q)=P$, since $Q$ is a $P$-primary ideal of $R$.
+Hence, either $sa\in Q\cap J$ or $sb\in P\cap J\subseteq P\cap rad(J)=rad(Q)\cap rad(J)=rad(Q\cap J)$.
+This proves that $Q\cap J$ is a $(P\cap rad(J))$-$S$-primary ideal of $R$.
+If an ideal $I$ of a ring $R$ admits an $S$-primary decomposition, then we prove in Remark 11 that it will admit a minimal $S$-primary decomposition.
 Following [3], let $E$ be a family of ideals of a ring $R$.
 An element $I\in E$ is said to be an $S$-maximal element of $E$ if there exists an $s\in S$ such that for each $J\in E$, if $I\subseteq J$, then $sJ\subseteq I$.
-Also an increasing sequence $(I_{i})_{i\in\mathbb{N}}$ of ideals of $R$ is called $S$-stationary if there exists a positive integer $k$ and $s\in S$ such that $sI_{n}\subseteq I_{k}$ for all $n\geq k$.
+Also an increasing sequence $(I_{i})_{i\in\mathbb{N}}$ of ideals of $R$ is called $S$-stationary if there exist a positive integer $k$ and $s\in S$ such that $sI_{n}\subseteq I_{k}$ for all $n\geq k$.
 The following theorem provides a connection between the concepts of $S$-irreducible ideals and $S$-primary ideals.
@@
 Suppose $Q$ is an $S$-irreducible ideal of $R$.
-Let $a,b\in R$ be such that $ab\in Q$ and $sb\notin Q$ for all $s\in S$.
-Our aim is to show that there exists $t\in S$ such that $ta\in rad(Q).$ Consider $A_{n}=\{x\in R\hskip 5.69046pt|\hskip 2.84544pta^{n}x\in Q\}$ for $n\in\mathbb{N}.$ Clearly each $A_{n}$ is an ideal of $R$ and $A_{1}\subseteq A_{2}\subseteq A_{3}\subseteq\cdots$ is an increasing chain of ideals of $R$.
-Since $R$ is $S$-Noetherian, then by [9, Theorem 2.3], this chain is $S$-stationary, i.e., there exist $k\in\mathbb{N}$ and $s\in S$ such that $sA_{n}\subseteq A_{k}$ for all $n\geq k$.
+Let $a,b\in R$ such that $ab\in Q$ and $sb\notin Q$ for all $s\in S$.
+Our aim is to show that there exists $t\in S$ such that $ta\in rad(Q)$.
+Consider $A_{n}=\{x\in R\hskip 5.69046pt|\hskip 2.84544pta^{n}x\in Q\}$ for $n\in\mathbb{N}.$ Clearly each $A_{n}$ is an ideal of $R$ and $A_{1}\subseteq A_{2}\subseteq A_{3}\subseteq\cdots$ is an increasing chain of ideals of $R$.
+Since $R$ is $S$-Noetherian, by [8, Theorem 2.3], this chain is $S$-stationary, i.e., there exist $k\in\mathbb{N}$ and $s\in S$ such that $sA_{n}\subseteq A_{k}$ for all $n\geq k$.
 Consider the two ideals $I=\textless a^{k}\textgreater+\hskip 2.84544ptQ$ and $J=\textless b\textgreater+\hskip 2.84544ptQ$ of $R$.
@@
 Theorem 9.
-(Existence of $S$-Primary Decomposition) Let $R$ be an $S$-Noetherian ring. Then every proper ideal (disjoint with $S$) of $R$ can be written as a finite intersection of $S$-primary ideals.
+(Existence of $S$-Primary Decomposition) Let $R$ be an $S$-Noetherian ring. Then every proper ideal of $R$ disjoint from $S$ can be written as a finite intersection of $S$-primary ideals.
 Proof.
-Let $E$ be a collection of ideals of $R$ which can not be written as a finite intersection of $S$-irreducible ideals.
+Let $E$ be the collection of ideals of $R$ which are disjoint from $S$ and can not be written as a finite intersection of $S$-primary ideals.
 We wish to show $E=\emptyset$.
-On contrary suppose $E\neq\emptyset$.
-Since $R$ is an $S$-Noetherian ring, by [9, Theorem 2.3], there exists an $S$-maximal element in $E$, say $I$.
-Evidently, $I$ is not an $S$-irreducible ideal, and so $I$ is not an irreducible ideal.
+On the contrary suppose $E\neq\emptyset$.
+Since $R$ is an $S$-Noetherian ring, by [8, Theorem 2.3], there exists an $S$-maximal element in $E$, say $I$.
+Evidently, $I$ is not an $S$-primary ideal, by Theorem 8, $I$ is not an $S$-irreducible ideal, and so $I$ is not an irreducible ideal.
 This implies that $I=J\cap K$ for some ideals $J$ and $K$ of $R$ with $I\neq J$ and $I\neq K$.
@@
 Now we claim that $J$, $K\notin E$.
-For this, suppose that $J$ and $K$ are in $E$, then there exist $s^{\prime},s^{\prime\prime}\in S$ such that $s^{\prime}J\subseteq I$ and $s^{\prime\prime}K\subseteq I$ and $I\subset J$, $I\subset K$ which is not possible.
+For this, if $J$ (respectively, $K$) belongs to $E$, then since $I$ is an $S$-maximal element of $E$ and $I\subset J$ (respectively, $I\subset K$), there exists $s^{\prime}$ (respectively, $s^{\prime\prime}$) from $S$ such that $s^{\prime}J\subseteq I$ (respectively, $s^{\prime\prime}K\subseteq I$).
+This is not possible, as $I$ is not $S$-irreducible.
 Therefore $J,K\notin E$.
-This implies that $J$ and $K$ can be written as a finite intersections of $S$-irreducible ideals.
-Consequently, $I$ can also be written as a finite intersection of $S$-irreducible ideals since $I=J\cap K$, a contradiction as $I\in E$.
-Thus $E=\emptyset$, i.e., every proper ideal of $R$ can be written as finite intersection of $S$-irreducible ideals.
-But by Theorem 10, an $S$-irreducible ideal is $S$-primary.
-Therefore every ideal of $R$ can be written as a finite intersection of $S$-primary ideals.
+This implies that $J$ and $K$ can be written as a finite intersection of $S$-primary ideals.
+Consequently, $I$ can also be written as a finite intersection of $S$-primary ideals since $I=J\cap K$, a contradiction as $I\in E$.
+Thus $E=\emptyset$, i.e., every proper ideal of $R$ disjoint from $S$ can be written as a finite intersection of $S$-primary ideals.
 ∎
-Recall that if $I$ is any ideal of $R$, the radical of $I$ is $rad(I)=\{x\in R\hskip 2.84544pt|\hskip 2.84544ptx^{n}\in I\hskip 2.84544ptfor\hskip 2.84544ptsome\hskip 2.84544ptn\geq 0\}$.
-If $I=rad(I)$, then $I$ is called radical ideal.
-As an application of this theorem, we provide the following result.
+Recall that if $I$ is any ideal of $R$, the radical of $I$ is $rad(I)=\{x\in R\hskip 2.84544pt|\hskip 2.84544ptx^{n}\in I\hskip 2.84544ptfor\hskip 2.84544ptsome\hskip 2.84544ptn\textgreater 0\}$.
+If $I=rad(I)$, then $I$ is called a radical ideal.
 Corollary 10.
-Let $R$ be an $S$-Noetherian ring. Then every radical ideal $I$ (disjoint with $S$) of $R$ is the intersection of finitely many $S$-prime ideals.
+Let $R$ be an $S$-Noetherian ring. Then every radical ideal $I$ (disjoint from $S$) of $R$ is the intersection of finitely many $S$-prime ideals.
 Proof.
 By Theorem 9, there exist finitely many $S$-primary ideals $Q_{1},Q_{2},\ldots,Q_{r}$ such that $I=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$.
-Also by [14, Proposition 2.5], $rad({Q_{i}})=P_{i}$ is $S$-prime for $i=1,2,\ldots,r$.
+Also by [13, Proposition 2.5], $rad({Q_{i}})=P_{i}$ is $S$-prime for $i=1,2,\ldots,r$.
 Consequently, $I=rad(I)=P_{1}\cap P_{2}\cap\cdots\cap P_{r}$.
@@
 Remark 11.
-Let $R$ be an $S$-Noetherian ring and $I$ be an ideal of $R$.
-Then by Theorem 11, $I$ has $S$-primary decomposition.
-Suppose $I=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$ is an $S$-primary decomposition of $I$.
-Since each $Q_{i}$ is $S$-primary, so by [16, Proposition 2.6], each $S^{-1}Q_{i}$ is primary.
-Consequently, $S^{-1}I=S^{-1}Q_{1}\cap S^{-1}Q_{2}\cap\cdots\cap S^{-1}Q_{r}$ is a primary decomposition of $S^{-1}I$
-To prove $S$-version of the $1st$ uniqueness theorem for $S$-Noetherian rings, we need the following lemmas:
-Lemma 14.
+Let $R$ be a ring, and let $S$ be a multiplicative set of $R$.
+Let $I$ be an ideal of $R$ such that $I\cap S=\emptyset$.
+Suppose that $I$ admits an $S$-primary decomposition.
+(In Theorem 9, we have shown that if $R$ is $S$-Noetherian, then any ideal $I$ of $R$ disjoint from $S$ admits an $S$-primary decomposition).
+Let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be an $S$-primary decomposition of $I$ with $Q_{i}$ is a $P_{i}$-$S$-primary ideal of $R$ for each $i\in\{1,2,\ldots,n\}$.
+Then, by [13, Proposition 2.7], $S^{-1}Q_{i}$ is $S^{-1}P_{i}$-primary, $S(I)=\bigcap\limits_{i=1}^{n}S(Q_{i})$ is a primary decomposition with $S(Q_{i})$ is a $S(P_{i})$-primary for each $i\in\{1,\ldots,n\}$.
+Let $1\leq i\leq n$.
+Let $s_{i}\in S$ be such that it satisfies the $S$-primary property of $Q_{i}$.
+Notice that $(P_{i}:s_{i})$ is a prime ideal of $R$ and $(Q_{i}:s_{i})$ is a $(P_{i}:s_{i})$-primary ideal of $R$.
+Let $s=\prod_{i=1}^{n}s_{i}$.
+Then $s\in S$.
+Observe that $(Q_{i}:s_{i})=(Q_{i}:s)$, $(P_{i}:s_{i})=(P_{i}:s)$, $S(Q_{i})=(Q_{i}:s)$, and $S(P_{i})=(P_{i}:s)$.
+From $I=\bigcap\limits_{i=1}^{n}Q_{i}$, it follows that $(I:s)=\bigcap\limits_{i=1}^{n}(Q_{i}:s)=\bigcap\limits_{i=1}^{n}S(Q_{i})=S(I)$.
+Let $k$ of the $S(P_{1}),\ldots,S(P_{n})$ be distinct.
+After a suitable rearrangement of $\{1,\ldots,n\}$, we can assume without loss of generality that $S(P_{1}),\ldots,S(P_{k})$ are distinct among $S(P_{1}),\ldots,S(P_{n})$.
+Let $A_{1}=\{j\in\{1,\ldots,n\}\mid S(Q_{j})~is~S(P_{1})$-primary}\}, …, $A_{k}=\{j\in\{1,\ldots,n\}\mid S(Q_{j})~is~S(P_{k})$-primary }\}.
+From the above discussion, it is evident that $1\in A_{1},\ldots,k\in A_{k}$, and $\{1,\ldots,n\}=\bigcup\limits_{t=1}^{k}A_{t}$.
+Let $1\leq t\leq k$.
+Notice that $\bigcap_{j\in A_{t}}S(Q_{j})$ is $S(P_{t})$-primary by [6, Lemma 4.3].
+It is convenient to denote $\bigcap_{j\in A_{t}}Q_{j}$ by $I^{\prime}_{t}$.
+Thus $(I:s)=S(I)=S(I^{\prime}_{1})\cap S(I^{\prime}_{2})\cap\cdots\cap S(I^{\prime}_{k})$ with $S(I^{\prime}_{t})$ is $S(P_{t})$-primary for each $t\in\{1,\ldots,k\}$ and $S(P_{1}),\ldots,S(P_{k})$ are distinct.
+After omitting those $S(I^{\prime}_{i})$ such that $S(I^{\prime}_{i})\supseteq\bigcap_{t\in\{1,\ldots,k\}\setminus\{i\}}S(I^{\prime}_{t})$ from the intersection, we can assume without loss of generality that $(I:s)=S(I)=\bigcap\limits_{t=1}^{k}S(I^{\prime}_{t})$ is a minimal primary decomposition of $S(I)$.
+Next, we claim that $I=(I:s)\cap(I+Rs)$.
+It is clear that $I\subseteq(I:s)\cap(I+Rs)$.
+Let $y\in(I:s)\cap(I+Rs)$.
+Then $ys\in I$ and $y=a+rs$ for some $r\in R$.
+This implies that $ys=as+rs^{2}$ and so, $rs^{2}\in I$.
+Hence, $r\in S(I)=(I:s)$.
+Therefore, $y=a+rs\in I$.
+This shows that $(I:s)\cap(I+Rs)\subseteq I$.
+Thus $I=(I:s)\cap(I+Rs)$ and hence, $I=(\bigcap\limits_{t=1}^{k}S(I^{\prime}_{t})\cap(I+Rs))=\bigcap\limits_{t=1}^{k}(S(I^{\prime}_{t})\cap(I+Rs))$.
+Let $1\leq t\leq k$.
+For convenience, let us denote $S(I^{\prime}_{t})\cap(I+Rs)$ by $Q^{\prime}_{t}$.
+As $S(I^{\prime}_{t})$ is $S(P_{t})$-primary with $S(I^{\prime}_{t})\cap S=\emptyset$ and $(I+Rs)\cap S\neq\emptyset$, we obtain from Proposition 7(2) that $Q^{\prime}_{t}$ is $(P_{t})\cap rad(I+Rs)-SS$-primary.
+Since $S(S(P_{t}))=S(P_{t})$, $S(rad(I+Rs))=R$, it follows that $S(S(P_{t})\cap rad(I+Rs))=S(P_{t})$.
+Notice that $S(Q^{\prime}_{t})=S(S(I^{\prime}_{t})\cap(I+Rs))=S(I^{\prime}_{t})$, as $S(S(I^{\prime}_{t}))=S(I^{\prime}_{t})$ and $S(I+Rs)=R$.
+Hence, for all distinct $i,j\in\{1,\ldots,k\}$, $S(S(P_{i})\cap rad(I+Rs))\neq S(S(P_{j})\cap rad(I+Rs))$ and for each $i$ with $1\leq i\leq t$, $S(Q^{\prime}_{i})\nsupseteq\bigcap_{t\in\{1,\ldots,k\}\setminus\{i\}}S(Q^{\prime}_{t})$.
+Therefore, $I=\bigcap\limits_{t=1}^{k}Q^{\prime}_{t}$ is a minimal $S$-primary decomposition.
+Recall from [2], let $R$ be a ring, $S\subseteq R$ a multiplicative set and $I$ an ideal of $R$ disjoint from $S$.
+Let $P$ be an $S$-prime ideal of $R$ such that $I\subseteq P$.
+Then $P$ is said to be a minimal $S$-prime ideal over $I$ if $P$ is minimal in the set of the $S$-prime ideals containing $I$.
+Also Ahmed [2, Remark 2] proves in $S$-Noetherian rings that the set of minimal $S$-prime ideals is finite if $S$ is a finite multiplicative set.
+A prime ideal is said to be a minimal prime ideal if it is a minimal prime ideal over the zero ideal.
+Emmy Noether showed that in a Noetherian ring, there are only finitely many minimal prime ideals over any given ideal [10, Theorem 88].
+A natural question arises:
+Question 2.1.
+Is the collection of minimal prime ideals in $S$-Noetherian rings finite for any multiplicative set $S$?
+The answer to the above question is negative. In the following example, we provide an $S$-Noetherian ring that has infinitely many minimal prime ideals.
+Example 12.
+Consider the ring $R=\dfrac{F[x_{1},x_{2},\ldots,x_{n},\ldots]}{(x_{i}x_{j};i\neq j,i,j\in\mathbb{N})}$, where $F$ is a field.
+Let $y_{i}={\overline{x}}_{i}$ be the image of $x_{i}$ under the canonical map.
+Consider the multiplicative set $S=\{y_{1}^{n}:n\in\mathbb{N}\cup\{0\}$} of $R$.
+Then $(y_{1})\subseteq(y_{1},y_{2})\subseteq\cdots\subseteq(y_{1},y_{2},\ldots,y_{n})\subseteq\cdots$ is an ascending chain of ideals of $R$ which is not stationary.
+Consequently, $R$ is not a Noetherian ring.
+Evidently, $R$ is an $S$-Noetherian ring (see [15, Example 4]).
+Note that if $P\subset R$ is a prime ideal, then there can be at most one $i$ such that $y_{i}\notin P$.
+It follows that if $P_{i}$ is the ideal generated by all the $y_{j}$ for $j\neq i$, each $P_{i}$ is prime since $R/P_{i}\cong F[x_{i}]$.
+Clearly, every $P_{i}$ is a minimal prime ideal of $R$.
+Thus $R$ has infinitely many minimal prime ideals.
+Question 2.2.
+Under what condition the set of minimal prime ideals is finite in $S$-Noetherian rings?
+We give an answer to the above question under a mild condition on the multiplicative set $S\subseteq R$ (see Theorem 14).
+Remark 13.
+Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $(0)=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary ideal of $R$.
+If $Z(R/P_{i})\cap\overline{S}_{i}=\emptyset$, where $\overline{S}_{i}=\{s+P_{i}\mid s\in S\}$ and $Z(R/P_{i})$ denotes the set of zero divisors of $R/P_{i}$ for all $i=1,2,\ldots,r$.
+Then each $P_{i}$ is a prime ideal of $R$.
+For this, we will show that $(P_{i}:s)=P_{i}$ for all $s\in S$ and $i=1,2,\ldots,r$.
+It is clear that $P_{i}\subseteq(P_{i}:s)$.
+Conversely, let $s\in S$ and $x\in(P_{i}:s)$, then $sx\in P_{i}$; so $(s+P_{i})(x+P_{i})=P_{i}$.
+Thus $x\in P_{i}$ since $Z(R/P_{i})\cap\overline{S}_{i}=\emptyset$.
+Now, if $P_{i}$ is a $S$-prime ideal of $R$, then by [2, Proposition 1], $(P_{i}:s_{i})$ is a prime ideal of $R$ for some $s_{i}\in S$.
+Since $P_{i}=(P_{i}:s_{i})$, $P_{i}$ is a prime ideal of $R$.
+Theorem 14.
+Let $R$ be an $S$-Noetherian ring, $S\subseteq R$ a multiplicative set, and $(0)=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary ideal of $R$.
+If $Z(R/P_{i})\cap\overline{S}_{i}=\emptyset$, then minimal prime ideals of $R$ are in the set $\{P_{i}\mid i=1,2,\ldots r\}$.
+Proof.
+Since $(0)=Q_{1}\cap Q_{2}\cap\cdots\cap Q_{r}$, $rad~(0)=P_{1}\cap P_{2}\cap\cdots\cap P_{r}$, where $P_{i}=rad(Q_{i})$ $(1\leq i\leq r)$ is an $S$-prime ideal of $R$ By Remark 13, each $P_{i}$ $(1\leq i\leq r)$ is a prime ideal of $R$.
+Let $P^{\prime}$ be any minimal prime ideal of $R$.
+Then $(0)\subseteq P^{\prime}$, and so $rad~(0)\subseteq rad(P^{\prime})=P^{\prime}$.
+This implies that $P_{1}\cap P_{2}\cap\cdots\cap P_{r}\subseteq P^{\prime}$.
+By [6, Proposition 1.11(ii)], there exists $j\in\{1,2,\ldots,r\}$ such that $P_{j}\subseteq P^{\prime}$.
+Since $P^{\prime}$ is a minimal prime ideal, $P^{\prime}=P_{j}$, as desired.
+∎
+Proposition 15.
 Let $Q_{1},Q_{2},\ldots,Q_{n}$ be ideals of $R$, and $P$ be an $S$-prime ideal containing $\bigcap\limits_{k=1}^{n}Q_{k}$.
@@
 Evidently, $x_{1}x_{2}\ldots x_{n}\in Q_{1}Q_{2}\ldots Q_{n}$.
-This implies that $(sx_{1})(sx_{2})\cdots(sx_{n})\in Q_{1}Q_{2}\ldots Q_{n}\subseteq\bigcap\limits_{k=1}^{n}Q_{k}\subseteq P$.
-Then there exist $s^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $s^{\prime}sx_{j}\in P$ since $P$ is an $S$-prime ideal, a contradiction.
+This implies that $x_{1}x_{2}\cdots x_{n}\in Q_{1}Q_{2}\ldots Q_{n}\subseteq\bigcap\limits_{k=1}^{n}Q_{k}\subseteq P$.
+Since $P$ is an $S$-prime ideal, by [2, Proposition 4], there exist $s^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $s^{\prime}x_{j}\in P$, a contradiction.
 Hence there exist $s\in S$ and $k\in\{1,2,\ldots,n\}$ such that $sQ_{k}\subseteq P$.
@@
 ∎
-Lemma 15.
-Let $Q$ be an $S$-primary ideal and $x$ an element of $R$ such that $sx\notin Q$ for all $s\in S$.
-Then $(Q:sx)$ is $S$-primary, and there exists $t\in S$ such that $t(rad(Q:sx))\subseteq rad(Q)\subseteq rad(Q:sx)$ for all $s\in S$.
+Now we prove $S$-version of the $1st$ uniqueness theorem:
+Theorem 16.
+$(\textbf{$S$-version of 1st $S$-uniqueness theorem})$.
+Let $R$ be a ring and let $S$ be a multiplicative set of $R$.
+Let $I$ be an ideal of $R$ which admits $S$-primary decomposition.
+Let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary for each $i\in\{1,2,\ldots,n\}$.
+Then the $S(P_{i})$ are precisely the prime ideals which occur in the set of ideals $S(rad(I:x))$ $(x\in R)$, and hence are independent of the particular $S$-decomposition of $I$.
 Proof.
-Suppose $(Q:sx)\cap S\neq\emptyset$ for some $s\in S$.
-Then there exists $t\in S$ such that $t\in(Q:sx)$ which imples that $tsx\in Q$, a contradiction.
-Thus $(Q:sx)\cap S=\emptyset$ for all $s\in S.$ Now let $s\in S$ and $yz\in(Q:sx)$ be such that $s^{\prime}y\notin rad(Q:sx)$ for all $s^{\prime}\in S$.
-Then $sxyz\in Q$.
-Since $Q$ is $S$-primary and $s^{\prime}y\notin rad(Q:sx)$ for all $s^{\prime}\in S$, there exists $s^{\prime\prime}\in S$ such that $s^{\prime\prime}sxz\in Q$.
-Consequently, $s^{\prime\prime}z\in(Q:sx)$.
-Thus $(Q:sx)$ is $S$-primary.
-Further, let $w\in rad(Q:sx)$.
-Then $w^{n}\in(Q:sx)$ for some $n\in\mathbb{N}$.
-This implies that $sxw^{n}\in Q$.
-Since $Q$ is an $S$-primary, there exists $t\in S$ such that $(tw^{n})^{k}\in Q$ for some $k$ as $tsx\notin Q$.
-This concludes that $(tw)^{nk}=t^{nk-k}(tw^{n})^{k}\in Q$ which implies that $tw\in rad(Q)$.
-Thus $t(rad(Q:sx))\subseteq rad(Q)$.
-Also $Q\subseteq(Q:sx)$ implies that $rad(Q)\subseteq rad(Q:sx)$, and hence $t(rad(Q:sx))\subseteq rad(Q)\subseteq rad(Q:sx)$.
-∎
-Theorem 16.
-[…]
-Proof.
-For any $x\in R$, we have $(I:x)=(\bigcap\limits_{i=1}^{n}Q_{i}:x)=\bigcap\limits_{i=1}^{n}(Q_{i}:x)$, and so $rad(I:x)=\bigcap\limits_{i=1}^{n}rad(Q_{i}:x)$.
-First we show that $rad(I:x)$ satisfies $sP_{i}\subseteq s^{\prime}(rad(I:x))\subseteq P_{i}$ for some $s,s^{\prime}\in S$ is $S$-prime.
-For this, let $\alpha\beta\in rad(I:x)$ such that $t\beta\notin rad(I:x)$ for all $t\in S$.
-This implies $s^{\prime}\alpha\beta\in s^{\prime}rad(I:x)\subseteq P_{i}$.
-Then there exists $s^{\prime\prime}\in S$ such that either $s^{\prime\prime}\alpha\in P_{i}$ or $s^{\prime}s^{\prime\prime}\beta\in P_{i}$ since $P_{i}$ is $S$-prime.
-If $s^{\prime}s^{\prime\prime}\beta\in P_{i}$, then $ss^{\prime}s^{\prime\prime}\beta\in sP_{i}\subseteq s^{\prime}(rad(I:x))\subseteq rad(I:x)$, a contradiction since $t\beta\notin rad(I:x)$ for all $t\in S$.
-Hence $s^{\prime\prime}\alpha\in P_{i}$, and so $ss^{\prime\prime}\alpha\in sP_{i}\subseteq s^{\prime}rad(I:x)\subseteq rad(I:x)$, i.e., $s^{\prime\prime\prime}\alpha\in rad(I:x)$, where $s^{\prime\prime\prime}=ss^{\prime\prime}\in S$.
-Thus $rad(I:x)$ is an $S$-prime ideal.
-Now by Lemma 14, there exist $t^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $t^{\prime}(rad(Q_{j}:x))\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$.
-Choose $x$ such that $tx\notin Q_{j}$ for all $t\in S$.
-Then by Lemma 15, there exists $t^{\prime\prime}\in S$ such that $t^{\prime\prime}(rad(Q_{j}:x))\subseteq rad(Q_{j})=P_{j}\subseteq rad(Q_{j}:x)$.
-This implies that $t^{\prime}t^{\prime\prime}(rad(Q_{j}:x))\subseteq t^{\prime}P_{j}\subseteq t^{\prime}(rad(Q_{j}:x))\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$.
-Thus we have $t^{\prime}P_{j}\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$, and so $t^{\prime\prime}t^{\prime}P_{j}\subseteq t^{\prime\prime}(rad(I:x))\subseteq t^{\prime\prime}(rad(Q_{j}:x))\subseteq rad(Q_{j})=P_{j}$.
-Put $s=t^{\prime\prime}t^{\prime}$ and $s^{\prime}=t^{\prime\prime}$.
-Then we have $sP_{j}\subseteq s^{\prime}(rad(I:x))\subseteq P_{j}$ for some $j$.
-Conversely, since decomposition is minimal, so for each $k$, $\bigcap\limits_{i\neq k}^{n}Q_{i}\nsubseteq Q_{k}$.
-This implies that there exists $x_{k}\in(\bigcap\limits_{i\neq k}^{n}Q_{i})\setminus Q_{k}$ for each $k$.
-Consequently, for each $k$, we have $(I:x_{k})=\bigcap\limits_{i=1}^{n}(Q_{i}:x_{k})=(Q_{k}:x_{k})$ since $(Q_{i}:x_{k})=R$ for all $i\neq k$ by [7, Lemma 4.4].
-Notice that $tx_{k}\notin Q_{k}$ for every $t\in S$, for if $tx_{k}\in Q_{k}$ for some $t\in S$, then $t\in(Q_{k}:x_{k})=(I:x_{k})\subseteq rad(I:x_{k})$, which is not possible since $rad(I:x_{k})$ is $S$-prime.
-So by Lemma 15, for each $k$, there exists $s_{k}\in S$ such that $s_{k}(rad(Q_{k}:x_{k}))\subseteq rad(Q_{k})=P_{k}$.
-Also $Q_{k}\subseteq(Q_{k}:x_{k})=(I:x_{k})$, so $P_{k}=rad(Q_{k})\subseteq rad(I:x_{k})$.
-Consequently, for each $k$, $s_{k}P_{k}\subseteq s_{k}(rad(I:x_{k}))=s_{k}(rad(Q_{k}:x_{k}))\subseteq P_{k}$.
-Put $s=s_{1}s_{2}\ldots s_{n}$.
-Then $sP_{k}\subseteq s(rad(I:x_{k}))\subseteq P_{k}$ for all $k=1,2,\ldots,n$.
+For any $x\in R$, we have $S^{-1}(I:x)=(\bigcap\limits_{i=1}^{n}S^{-1}Q_{i}:_{S^{-1}R}\frac{x}{1})=\bigcap\limits_{i=1}^{n}(S^{-1}Q_{i}:_{S^{-1}R}\frac{x}{1})$, and so $S(I:x)=\bigcap\limits_{i=1}^{n}(S(Q_{i}):x)$.
+Evidently, $S(rad(I:x))=\bigcap\limits_{i=1}^{n}rad(S(Q_{i}):x)$.
+Notice that each $S(P_{i})$ is a prime ideal since $S^{-1}P_{i}$ is a prime ideal and contraction of a prime ideal is a prime ideal.
+Now we show that each $S(Q_{i})$ is $S(P_{i})$-primary.
+For this, let $x,~y\in R$ such that $xy\in S(Q_{i})$.
+Then $\dfrac{xy}{1}\in S^{-1}Q_{i}$, $\dfrac{xy}{1}=\dfrac{a}{s}$ for some $a\in Q_{i}$ and $s\in S$.
+Consequently, there exists $u\in S$ such that $usxy\in Q_{i}$.
+Since $Q_{i}$ is $P_{i}$-$S$-primary, there exists $t\in S$ such that either $tusx\in Q_{i}$ or $ty\in P_{i}$.
+It follows that either $\dfrac{x}{1}=\dfrac{tusx}{tus}\in S^{-1}Q_{i}$ or $\dfrac{y}{1}=\dfrac{ty}{t}\in S^{-1}P_{i}$.
+This implies that either $x\in S(Q_{i})$ or $y\in S(P_{i})$, and rad $S(Q_{i})=S(rad(Q_{i}))=S(P_{i})$.
+Therefore $S(Q_{i})$ is $S(P_{i})$-primary.
+Then $S(rad(I:x))=\bigcap\limits_{i=1}^{n}rad(S(Q_{i}):x)=\bigcap\limits_{x\notin S(Q_{i})}S(P_{i})$, by [6, Lemma 4.4].
+Suppose $S(rad(I:x))$ is prime; then we have $S(rad(I:x))=S(P_{j})$ for some $j$, by [6, Proposition 1.11].
+Hence every prime ideal of the form $S(rad(I:x))$ is one of the $S(P_{j})$.
+Conversely, for each $i$ there exists $x_{i}\in\left(\bigcap_{j\in\{1,2,\ldots,n\}\setminus\{i\}}S(Q_{j})\right)\setminus S(Q_{i})$ since the $S$-decomposition is minimal.
+Consequently, for each $i$, we have $S(I:x_{i})=\bigcap\limits_{i=1}^{n}(S(Q_{i}):x_{i})=(S(Q_{i}):x_{i})$ since $(S(Q_{j}):x_{i})=R$ for all $j\neq i$, by [6, Lemma 4.4].
+Then $S(rad(I:x_{i}))=S(P_{i})$, by [6, Lemma 4.4].
 This completes the proof.
 ∎
-For any ideal $I$ of a ring $R$ and any multiplicative set $S\subseteq R$, the contraction of ideal $S^{-1}I$ is denoted by $S(I)=\{a\in R\hskip 5.69046pt|\hskip 5.69046pt\dfrac{a}{1}\in S^{-1}I\}$.
-Notice that $I\subseteq S(I)$.
 To prove the second uniqueness theorem for the $S$-Noetherian rings, we need the following results:
 Lemma 17.
-Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ be an $S$-$P$-primary ideal of $R$.
-If $S\cap P=\emptyset$, then $S^{-1}Q$ is $S^{-1}P$-primary and its contraction in $R$ satisfies $tS(Q)\subseteq Q\subseteq S(Q)$ for some $t\in S$.
+Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ be an $P$-$S$-primary ideal of $R$.
+Then $S^{-1}Q$ is $S^{-1}P$-primary and its contraction in $R$ satisfies $tS(Q)\subseteq Q\subseteq S(Q)$ for some $t\in S$.
 Proof.
-Clearly $S^{-1}Q\neq S^{-1}R$, for $S\cap P=\emptyset$ and $P=rad(Q)$.
+Clearly, $S^{-1}Q\neq S^{-1}R$ since $Q\cap S=\emptyset$ and $P=rad(Q)$.
 Let $x\in S(Q)$.
@@
 Consequently, there exists $s^{\prime}\in S$ such that $(xs-a)s^{\prime}=0$; whence $ss^{\prime}x=s^{\prime}a\in Q$.
-Put $t=ss^{\prime}$.
-Then $tx\in Q$, and so $tS(Q)\subseteq Q\subseteq S(Q)$.
-Now since $Q$ is $S$-primary, by [16, Proposition 2.6], $S^{-1}Q$ is a primary ideal of $S^{-1}R$ and $rad(S^{-1}Q)=S^{-1}(rad(Q))=S^{-1}P$.
+As $Q$ is $S$-primary, there exists $t\in S$ such that either $ss^{\prime}t\in rad(Q)=P$ or $tx\in Q$.
+This implies that $tx\in Q$ since $P\cap S=\emptyset$, and so $tS(Q)\subseteq Q\subseteq S(Q)$.
+Now since $Q$ is $S$-primary, by [13, Proposition 2.7], $S^{-1}Q$ is a primary ideal of $S^{-1}R$ and $rad(S^{-1}Q)=S^{-1}(rad(Q))=S^{-1}P$.
 Also since $P$ is an $S$-prime ideal of $R$, by [2, Remark $1$], $S^{-1}P$ is a prime ideal of $S^{-1}R$.
 Thus $S^{-1}Q$ is an $S^{-1}P$-primary ideal of $S^{-1}R$.
+∎
 Lemma 18.
-Let $R$ be an $S$-Noetherian ring, and $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $S$-$P_{i}$-primary.
-If $S\cap\left(\bigcap\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ for some $m$.
-Then $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$ and $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S(I)$ for some $t\in S$.
+Let $R$ be a ring and let $S$ be a multiplicative set of $R$.
+Let $I$ be an ideal of $R$ such that $I$ admits an $S$-primary decomposition.
+Let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition where $Q_{i}$ is $P_{i}$-$S$-primary for each $i\in\{1,2,\ldots,n\}$.
+Let $S^{\prime}$ be a multiplicative set of $R$ such that $S\subseteq S^{\prime}$.
+If $Q_{1},Q_{2},\ldots,Q_{m}$ are such that $Q_{i}\cap S^{\prime}=\emptyset$ for each $i\in\{1,\ldots,m\}$ and $Q_{i}\cap S^{\prime}\neq\emptyset$ for each $i\in\{m+1,\ldots,n\}$, then $S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$ is a primary decomposition and there exists $t\in S$ such that $tS^{\prime}(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S^{\prime}(I)$.
 Proof.
-By [7, Proposition 3.11(v)], $S^{-1}I=\bigcap\limits_{i=1}^{n}S^{-1}Q_{i}$.
-Since $S\cap Q_{i}\neq\emptyset$ for $i=m+1,m+2,\ldots,n$, so by [7, Proposition 4.8(i)], $S^{-1}Q_{i}=S^{-1}R$ for $i=m+1,m+2,\ldots,n$.
-Consequently, $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$.
-Also by Lemma 17, $S^{-1}Q_{i}$ is $S^{-1}P_{i}$-primary for $i=1,2,\ldots,m$.
-Since $P_{i}$ are distinct, $S^{-1}P_{i}$ are also distinct, and so $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$ is a minimal $S$-primary decomposition.
-Next, let $x\in S(I)$.
-Then $\dfrac{x}{1}\in S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$.
-Write $\dfrac{x}{1}=\dfrac{a_{i}}{s_{i}}$ for some $a_{i}\in Q_{i}$ and $s_{i}\in S$.
-This implies that $(xs_{i}-a_{i})s=0$ for some $s\in S$; whence $s_{i}sx=sa_{i}\in Q_{i}$ for $i=1,2,\ldots,m$.
-Put $t=ss_{1}s_{2}\ldots s_{m}$.
-Then $tx\in\bigcap\limits_{i=1}^{m}Q_{i}$, and hence $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}$.
-Also by [7, Proposition 1.17(i)], $\bigcap\limits_{i=1}^{m}Q_{i}\subseteq\bigcap\limits_{i=1}^{m}S(Q_{i})=S(I)$.
-Hence $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S(I)$.
+By [6, Proposition 3.11(v)], $S^{\prime-1}I=\bigcap\limits_{i=1}^{n}S^{\prime-1}Q_{i}$.
+Since $S^{\prime}\cap Q_{i}\neq\emptyset$ for $i=m+1,m+2,\ldots,n$, by [6, Proposition 4.8(i)], $S^{\prime-1}Q_{i}=S^{\prime-1}R$ for $i=m+1,m+2,\ldots,n$.
+Consequently, $S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$.
+As $Q_{i}$ is $P_{i}$-$S$-primary and $S\subseteq S^{\prime}$, it follows that $Q_{i}$ is $P_{i}$-$S^{\prime}$-primary for each $i=1,\ldots,m$.
+Also by Lemma 17, $S^{\prime-1}Q_{i}$ is $S^{\prime-1}P_{i}$-primary for $i=1,2,\ldots,m$.
+Thus $S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$ is a primary decomposition.
+Next, let $x\in S^{\prime}(I)$.
+Then $\dfrac{x}{1}\in S^{\prime-1}I=\bigcap\limits_{i=1}^{m}S^{\prime-1}Q_{i}$.
+Write $\dfrac{x}{1}=\dfrac{a_{i}}{s_{i}}$ for some $a_{i}\in Q_{i}$ and $s_{i}\in S^{\prime}$.
+This implies that $(xs_{i}-a_{i})s=0$ for some $s\in S^{\prime}$; whence $s_{i}sx=sa_{i}\in Q_{i}$ for $i=1,2,\ldots,m$.
+Then there exists $t_{i}\in S$ such that $t_{i}x\in Q_{i}$ since $Q_{i}$ is $P_{i}$-$S$-primary and $P_{i}\cap S^{\prime}=\emptyset$ for all $i=1,2,\ldots,m$.
+Put $t=t_{1}t_{2}\ldots t_{m}$.
+Then $tx\in\bigcap\limits_{i=1}^{m}Q_{i}$, and hence $tS^{\prime}(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S^{\prime}(\bigcap\limits_{i=1}^{m}Q_{i})=\bigcap\limits_{i=1}^{m}S^{\prime}(Q_{i})=S^{\prime}(I)$.
+Hence $tS^{\prime}(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S^{\prime}(I)$.
 ∎
-For an ideal $I$ of $R$, let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $S$-$P_{i}$-primary.
+For an ideal $I$ of $R$, let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $P_{i}$-$S$-primary.
 Then the $S$-prime ideals $P_{i}$ are said to belong to $I$.
@@
 ($S$-version of 2nd uniqueness theorem).
-Let $R$ be an $S$-Noetherian ring and $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition of $I$, where $Q_{i}$ is $S$-$P_{i}$-primary.
+Let $I$ be an ideal of $R$ which admits $S$-primary decomposition, that is, $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition of $I$, where $Q_{i}$ is $P_{i}$-$S$-primary.
 If $\{P_{{1}},\ldots,P_{{m}}\}$ is a set of isolated $S$-prime ideals of $I$ for some $m$ $(1\leq m\leq n)$, then $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition.
 Proof.
-Since each $P_{i}$ is $S$-prime, so by [2, Proposition 1], there exist $s_{i}\in S$ such that $(P_{i}:s_{i})$ is a prime ideal for $i=1,2,\ldots,m$.
-Consider a multiplicative set $S^{\prime}=R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$ of $R$.
+Since each $P_{i}$ is $S$-prime, by [2, Proposition 1], there exists $s_{i}\in S$ such that $(P_{i}:s_{i})$ is a prime ideal for $i=1,2,\ldots,m$.
+Consider the multiplicative set $S^{\prime}=R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$ of $R$.
 Evidently, $S^{\prime}\cap P_{i}=\emptyset$ for $i=1,2,\ldots m$ since $P_{i}\subseteq(P_{i}:s_{i})$.
 We claim that $S^{\prime}\cap P_{k}\neq\emptyset$ for $k=m+1,m+2,\ldots,n$.
-On contary $S^{\prime}\cap P_{k}=\emptyset$ for some $k$.
-This implies that $P_{k}\subseteq\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$, so by [7, Proposition 1.11(i)], $P_{k}\subseteq(P_{j}:s_{j})$ for some $j$ $(1\leq j\leq m)$.
+On the contrary $S^{\prime}\cap P_{k}=\emptyset$ for some $k$.
+This implies that $P_{k}\subseteq\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$, by [6, Proposition 1.11(i)], $P_{k}\subseteq(P_{j}:s_{j})$ for some $j$ $(1\leq j\leq m)$.
 Consequently, $s_{j}P_{k}\subseteq P_{j}$, a contradiction since $P_{j}$ is isolated $S$-prime.
-This concludes that $S^{\prime}\cap\left(\bigcap\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S^{\prime}\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ since $P_{i}=rad(Q_{i})$ for all $i=1,2,\ldots,n.$ Now since each $P_{i}$ is $S$-prime, $S\cap P_{i}=\emptyset$ for all $i=1,2,\ldots,m$.
+This concludes that $S^{\prime}\cap\left(\bigcup\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S^{\prime}\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ since $P_{i}=rad(Q_{i})$ for all $i=1,2,\ldots,n.$ Now since each $P_{i}$ is $S$-prime, $S\cap P_{i}=\emptyset$ for all $i=1,2,\ldots,m$.
 This implies that $S\cap(P_{i}:s_{i})=\emptyset$ for all $i=1,2,\ldots,m$.
-For, if $S\cap(P_{i}:s_{i})\neq\emptyset$ for some $i$, then there exits $s^{\prime}\in S$ such that $s^{\prime}s_{i}\in P_{i}$, a contradiction.
+For, if $S\cap(P_{i}:s_{i})\neq\emptyset$ for some $i$, then there exists $s^{\prime}\in S$ such that $s^{\prime}s_{i}\in P_{i}$, a contradiction.
 Consequently, $S\cap(\bigcup\limits_{i=1}^{m}(P_{i}:s_{i}))=\emptyset$, and so $S\subseteq R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})=S^{\prime}$.
-Then by [5, Proposition 2(d)], $R$ is $S^{\prime}$-Noetherian.
-Also since each $Q_{i}$ is $S$-primary and $S^{\prime}\cap Q_{i}=\emptyset$ for all $i$ $(1\leq i\leq m)$, so $Q_{i}$ is $S^{\prime}$-primary for all $i=1,2,\ldots,m.$ Thus by Lemma 18, we have $tS(I)\subseteq\bigcap\limits_{j=1}^{m}Q_{j}\subseteq S(I)$ for some $t\in S$.
-Hence $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition.
+Also since each $Q_{i}$ is $S$-primary and $S^{\prime}\cap Q_{i}=\emptyset$ for all $i$ $(1\leq i\leq m)$, $Q_{i}$ is $S^{\prime}$-primary for all $i=1,2,\ldots,m.$ Thus by Lemma 18, we have $tS^{\prime}(I)\subseteq\bigcap\limits_{j=1}^{m}Q_{j}\subseteq S^{\prime}(I)$ for some $t\in S$.
+From above it is clear that $Q_{{1}}\cap\cdots\cap Q_{{m}}$ depends only $I$, and hence $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition.
 ∎
+Acknowledgement: Authors sincerely thank the anonymous referee for thorough review and very useful comments and suggestions that helped to improve this paper.